### v2

# Constraints on the primordial curvature perturbations on small scales

Abstract
The power spectrum of the primordial curvature perturbation $\mathcal{P}_{\mathcal{R}}$ has been measured with high precision on large scales $10^{-4}\lesssim k\lesssim 3~{}\rm Mpc^{-1}$, basing on the observations of cosmic microwave background, Lyman-$\alpha$ forest and large scale structure. On small scales $3\lesssim k\lesssim 10^{23}~{}\rm Mpc^{-1}$, the constrains are mainly from the studies on the primordial black holes (PBHs). Specifically, on small scales $10^{17}\lesssim k\lesssim 10^{23}~{}{\rm Mpc^{-1}}$, the limits arise from studies on the lightest supersymmetric particles produced by PBHs radiation and the stable Planck-mass relics after its evaporation. It has been demonstrated that the big bang nucleosynthesis can be used to constrain the initial fraction of PBHs with masses $10^{9}\lesssim M_{\rm PBH}\lesssim 10^{13}~{}{\rm g}$, corresponding to the scales $10^{16}\lesssim k\lesssim 10^{18}~{}{\rm Mpc^{-1}}$. Recently, on one hand, it is found that the evaporation of light PBHs ($M_{\rm PBH}\lesssim 10^{9}\rm g$) can modify the expansion rate of the Universe and the baryon-to-photon ratio, resulting in the influences on the primordial abundance of light nuclei. On the other hand, it has been proposed that the ‘memory burden’ effect can slow down the mass loss rate of black hole (BH), leading to the existence of light PBHs by now. Based on the recent theoretical research process of BH and the limits on the (initial) mass fraction of light PBHs with masses $10^{4}\lesssim M_{\rm PBH}\lesssim 10^{10}~{}\rm g$, we derive new constraints on $\mathcal{P}_{\mathcal{R}}$ on small scales $1.5\times 10^{18}\lesssim k\lesssim 2.5\times 10^{21}~{}\rm Mpc^{-1}$, which are rarely studied in previous literature.
I Introduction
It is well known that the cosmic structures come from early density perturbations with a amplitude of $\delta\rho/\rho\sim 10^{-5}$. The primordial curvature (density) perturbation power spectrum $\mathcal{P}_{\mathcal{R}}$ ($\mathcal{P}_{\delta}$) has been measured precisely on large scales, $\mathcal{P}_{\mathcal{R}}\sim 10^{-8}$ for $10^{-4}\lesssim k\lesssim 3~{}\rm Mpc^{-1}$, basing on the studies of cosmic microwave background (CMB), Lyman-$\alpha$ forest and large scale structure Hlozek et al. (2012); Bird et al. (2011); Tinker et al. (2012); Ragavendra et al. (2024). On small scales $3\lesssim k\lesssim 10^{23}~{}\rm Mpc^{-1}$, the constraints on $\mathcal{P}_{\mathcal{R}}$ mainly come from the studies of primordial black holes (PBHs), which are proposed to be formed in very early universe via the collapse of large density perturbation with a amplitude of $\delta\rho/\rho\gtrsim 0.3$, and the upper limits are $\mathcal{P}_{\mathcal{R}}\lesssim 10^{-2}$ Josan et al. (2009); Yang (2019); Dalianis (2019); Sato-Polito et al. (2019); Yang (2020a); Mack and Wesley (2008); Caravano et al. (2024a, b). It has been proposed that on scales $3\lesssim k\lesssim 10^{7}~{}\rm Mpc^{-1}$, the upper limits on $\mathcal{P}_{\mathcal{R}}$ can be obtained by studying ultracompact minihalos (UCMHs) Bringmann et al. (2012); Li et al. (2012); Clark et al. (2016); Yang et al. (2013); Josan and Green (2010), which could be formed via the collapse of early density perturbation with a amplitude of $10^{-3}\lesssim\delta\rho/\rho\lesssim 0.3$ Ricotti and Gould (2009). The limit from UCMHs is $\mathcal{P}_{\mathcal{R}}\lesssim 10^{-7}$, which is much stronger than that of PBHs, but depends on the nature of dark matter particle Bringmann et al. (2012); Li et al. (2012); Clark et al. (2016); Yang et al. (2013); Josan and Green (2010).
The life time of a black hole (BH) with a mass $M_{\rm BH}$ is $\tau\sim 10^{64}(M_{\rm BH}/M_{\odot})^{3}\rm yr$. Therefore, PBHs with mass $M_{\rm pbh}<10^{15}\rm g$ have evaporated completely by now. The energy released from light PBHs has influences on the big bang nucleosynthesis (BBN) ($10^{9}\lesssim M_{\rm pbh}\lesssim 10^{13.5}~{}\rm g$) and CMB ($10^{13.5}\lesssim M_{\rm pbh}\lesssim 10^{15}~{}\rm g$Ricotti and Gould (2009); Zhang et al. (2007). The constraints on the initial mass fraction of PBHs $\beta_{\rm PBH}$ and then on $\mathcal{P}_{\mathcal{R}}$, corresponding to the scales $10^{15}\lesssim k\lesssim 10^{18}~{}\rm Mpc^{-1}$, can be obtained using the related observations. PBHs with mass $M_{\rm pbh}<10^{9}\rm g$ have a lifetime shorter than $\sim 0.01\rm s$, and they are free from BBN constrainsCarr et al. (2010). Recently, the authors of Boccia et al. (2024) investigated this issue again. They found that the evaporation of PBHs with masses $10^{8}\lesssim M_{\rm pbh}\lesssim 10^{9}~{}\rm g$ can modify the expansion rate of the Universe and the baryon-to-photon ratio, resulting in the influences on the primordial abundance of light nuclei. Therefore, the observed abundance of helium (${}^{4}{\rm He}$) and deuterium ($\rm D$) can be used to derive the upper limits on the initial mass fraction of PBHs for above mass range Boccia et al. (2024).
As mentioned above, light PBHs with mass $M_{\rm PBH}\lesssim 10^{15}\rm g$ do not survive by now due to Hawking radiation. However, the back-reaction of emission on the quantum state of the BH is ignored in Hawking radiation scenario, which is important when the energy of the emitted quanta is comparable to the total energy of BH Preskill (1992); Dvali et al. (2020). Recently, the ‘memory burden’ effect caused by the back-reaction is proposed, which can slow down the mass loss rate of BH Dvali (2018); Dvali et al. (2020, 2024); Thoss et al. (2024). Therefore, PBHs with mass $M_{\rm pbh}\lesssim 10^{15}\rm g$ can survive and evaporate by now. For instance, at present, a memory-burdened PBH with mass $M_{\rm PBH}\sim 10^{8}\rm g$ can still emit high energy photons and neutrinos Chianese et al. (2024). In general, the ‘memory burden’ effect works, e.g., when the mass of the BH becomes half its initial mass due to Hawking radiation. Therefore, two memory-burdened PBHs could merger to form a new BH, which can then emit high energy particles [^1] Here we have assumed that the mass of a memory-burdened PBH does not change significantly due to radiation.. The authors of Zantedeschi and Visinelli (2024) investigated this issue and obtained the constrains on the mass fraction of PBHs for the mass range of $10^{4}\lesssim M_{\rm pbh}\lesssim 10^{9}~{}\rm g$. Based on these new constraints on the (initial) mass fraction of light PBHs, in this work, we will derive new constraints on the power spectrum of primordial curvature perturbation on small scales.
This paper is organized as follows. In Sec. II we first briefly view the initial mass fraction of PBHs in theory. Then we review the new constraints on the (initial) mass fraction of PBHs based on the related observations. Finally, we derive new constraints on the primordial curvature perturbation power spectrum on small scales. The summary and conclusions are given in Sec. III.
II constraints on light PBHs and primordial curvature perturbation power spectrum on small scales
II.1 The initial mass fraction of PBHs
Many mechanisms have been proposed for the formation of PBHs, and the most popular of which is that PBHs are formed via the collapse of early large density perturbations when they enter the particle horizon see, e.g., Ref. Carr et al. (2010) and references therein. In case of Gaussian distribution of early density perturbations, the probability distribution of the smoothed density contrast can be written as Carr et al. (2010); Sato-Polito et al. (2019)
$P(\delta(R))=\frac{1}{\sqrt{2\pi}\delta(R)}{\rm exp}\left(-\frac{\delta^{2}(R)}{2\sigma^{2}(R)}\right),$ (1)
where the mass variance $\sigma^{2}(R)$ is in the form of
$\sigma^{2}(R)=\int^{\infty}_{0}W^{2}(kR)\mathcal{P}_{\delta}(k,t)\frac{dk}{k}$ (2)
with $W(kR)={\rm exp}(-k^{2}R^{2}/2)$. $\mathcal{P}_{\delta}$ is the primordial density perturbation power spectrum. In Press-Schechter throry, the initial mass fraction of PBHs can be written as Carr et al. (2010)
[…]
$\simeq$ ${\rm erfc}\left(\frac{\delta_{c}}{\sqrt{2\pi}\sigma(R)}\right),$ (3)
where $\delta_{c}$ is the critical density contrast required for the formation of PBHs, which depends on many factors. A simple argument implies $\delta_{c}\sim 0.3$ Carr et al. (2010), corresponding to the value of equation of state of matter at the radiation dominated epoch. Previous numerical simulations have found that the value of $\delta_{c}$ is in the range of $0.42\lesssim\delta_{c}\lesssim 0.66$ Polnarev and Musco (2007), depending on the curvature profile. Here we adopt $\delta_{c}=0.42$ for our calculations.
The primordial density perturbation power spectrum $\mathcal{P}_{\delta}$ is related to the primordial curvature perturbation power spectrum $\mathcal{P}_{\mathcal{R}}$ as Josan et al. (2009); Bringmann et al. (2012); Sato-Polito et al. (2019)
$\mathcal{P}_{\delta}=\frac{4(1+w)^{2}}{(5+3w)^{2}}\left(\frac{k}{aH}\right)^{4}\mathcal{P}_{\mathcal{R}},$ (4)
where $w$ is $1/3$ in the radiation dominated epoch.
In view of the conservation of the comoving number of PBHs, the initial mass fraction of PBHs, $\beta(M_{\rm PBH})$, is related to the present abundance $f_{\rm PBH}$ as Josan et al. (2009); Carr et al. (2010)
$\beta(M_{\rm PBH})=1.6\times 10^{-25}f_{\rm PBH}\left(M_{\rm PBH}/{\rm g}\right)^{1/2},$ (5)
where $f_{\rm PBH}=\Omega_{\rm PBH}/\Omega_{\rm DM}$ [^2] For deriving Eq. (5), we have adopted that 1) the fraction of the horizon mass collapsing into PBHs is $\gamma=0.2$; 2) the number of relativistic degrees of freedom at the formation time is $\rm g_{\star i}\approx 100$ Carr et al. (2010).. The (initial) mass fraction of PBHs has been constrained by many different ways see, e.g., Refs Carr et al. (2010); Su et al. (2024); Cang et al. (2021); Tashiro and Kadota (2021); Mena et al. (2019); Lu and Wu (2019); Cai et al. (2020); Chen et al. (2016); Inoue and Kusenko (2017); Ali-Haïmoud and Kamionkowski (2017); Clark et al. (2018); De Luca et al. (2021); Yang (2020b); Khlopov (2010) and references therein. Here we will review the new constraints on light PBHs with masses $10^{4}\lesssim M_{\rm PBH}\lesssim 10^{10}~{}\rm g$, which are rarely addressed in previous works.
II.2 Constraints on the (initial) mass fraction of light PBHs
New constrains from BBN. The influences of light PBHs with masses $10^{9}\lesssim M_{\rm pbh}\lesssim 10^{13.5}{\rm g}$ on the BBN have been widely investigated in previous works, which mainly focus on the effects of the high energy particles from evaporating PBHs on nuclear reactions Carr et al. (2010). Recently, the authors of Boccia et al. (2024) found that light PBHs with masses $10^{8}\lesssim M_{\rm pbh}\lesssim 10^{9}~{}{\rm g}$ have also effects on the BBN. The key point is that the expansion rate of the Universe and the baryon-to-photon ratio are changed due to these evaporating light PBHs, leading to the changes of the primordial abundance of light nuclei yields. Based on the observed abundance of helium ($\rm{}^{4}{He}$) and deuterium ($\rm D$), the authors of Boccia et al. (2024) obtained the constrains on the initial mass fraction of light PBHs in the mass range of $10^{8}\lesssim M_{\rm pbh}\lesssim 10^{9}~{}{\rm g}$, which has not been covered in previous works. The new constraints on the initial mass fraction of PBH $\beta_{\rm PBH}$ from BBN are shown in Fig. 1 (green shaded area with right $y$ axis) [^3] Since $\beta_{\rm PBH}$ is weakly constrained from the helium abundance ($\rm{}^{4}{He}$), here we show only the constraints from the deuterium abundance ($\rm D$)..
New constrains with ‘memory burden’ effect. According to the Hawking radiation theory, the lifetime of light PBHs with mass $M_{\rm PBH}<10^{15}\rm g$ is shorter than the current age of the Universe, which implies that they have evaporated completely by now. However, the back-reaction of emission on the quantum state of the BH is not included in Hawking radiation scenario, which is important when the energy of the emitted quanta is comparable to the total energy of BH Dvali (2018); Dvali et al. (2020, 2024); Thoss et al. (2024).
Recently, the ‘memory burden’ effect caused by the back-reaction has been proposed Dvali (2018); Dvali et al. (2020, 2024); Thoss et al. (2024). In this scenario, after losing, e.g., half mass of a PBH, the mass loss rate of a PBH, $dM_{\rm PBH}/dt$, is slowed down as $dM^{{}^{\prime}}_{\rm PBH}/dt=(4\pi GM^{2}_{\rm PBH})^{-k}dM_{\rm PBH}/dt$ with $k>0$. Therefore, light PBHs with mass $M_{\rm PBH}<10^{15}\rm g$ can still exist by now, and these memory-burdened PBHs also emit high energy particles, e.g., neutrinos and photons Chianese et al. (2024); Thoss et al. (2024). On the other hand, two memory-burdened PBHs could merger to form a new BH with a mass approximately equal to the sum of two original BHs, and the new formed BH can emit high energy particles via the Hawking radiation Zantedeschi and Visinelli (2024). Based on the observations of high energy neutrinos and gamma-rays, the authors of Zantedeschi and Visinelli (2024); Thoss et al. (2024); Chianese et al. (2024) derived the constrains on the fraction of light PBHs $f_{\rm PBH}$, and the results are shown in Fig. 1 (blue shaded area from gamma-rays, red shaded area from neutrinos and black shaded area from neutrinos for mergered PBHs).
Figure 1: New constraints on the (initial) mass fraction of light PBHs for the mass range of $3.3\times 10^{3}\lesssim M_{\rm pbh}\lesssim 8.5\times 10^{9}~{}\rm g$: 1) the initial mass fraction ($\beta_{\rm PBH}$) from BBN (the abundance of deuterium, green shaded area with right $y$ axis) Boccia et al. (2024); 2) the present mass fraction ($f_{\rm PBH}$) from gamma-rays for the memory-burdened PBHs (blue shaded area) Thoss et al. (2024); 3) the present mass fraction ($f_{\rm PBH}$) from high energy neutrinos for the memory-burdened PBHs (for the 7-year IceCube-EHE data, red shaded area) Chianese et al. (2024); 4) the present mass fraction ($f_{\rm PBH}$) from high energy neutrinos for the mergered memory-burdened PBHs (for the 7-year IceCube data with energy above 60 TeV, black shaded area) Zantedeschi and Visinelli (2024). All of shaded areas are excluded by the related observations.
II.3 Constraints on the primordial curvature perturbation power spectrum on small scales
The new constraints on the (initial) mass fraction of light PBHs, as shown in Fig. 1, can be used to derive the limits on the primordial curvature perturbation power spectrum on small scales. The final constrains on $\mathcal{P}_{\mathcal{R}}$ are shown in Fig. 2, corresponding to the scale range of $1.5\times 10^{18}\lesssim k\lesssim 2.5\times 10^{21}~{}\rm Mpc^{-1}$. On scales $10^{18}\lesssim k\lesssim 3\times 10^{20}~{}\rm Mpc^{-1}$, the upper limits are $\mathcal{P}_{\mathcal{R}}\lesssim 10^{-1.7}$, which are from the high energy neutrinos observations for mergered memory-burdened PBHs. The non-mergered memory-burdened PBHs provide the strongest limits on scales $3.5\times 10^{20}\lesssim k\lesssim 2.0\times 10^{21}~{}\rm Mpc^{-1}$, and the constraints from high energy neutrinos are comparable with that from gamma-rays. The constrains from the BBN, corresponding to the scales $4.6\times 10^{18}\lesssim k\lesssim 1.3\times 10^{19}~{}\rm Mpc^{-1}$, are weaker than that from memory-burdened PBHs.
Figure 2: Constraints on the power spectrum of primordial curvature perturbation ($\mathcal{P}_{\mathcal{R}}$) on small scales $1.5\times 10^{18}~{}{\rm Mpc^{-1}}\lesssim k\lesssim 2.5\times 10^{21}~{}\rm Mpc^{-1}$. The line style is the same as in Fig. 1.
III Summary and conclusions
Constraints on the primordial curvature perturbation power spectrum on small scales are mainly from the studies of PBHs. Specially, in previous works, on small scales $10^{17}\lesssim k\lesssim 10^{23}~{}{\rm Mpc^{-1}}$, the limits on $\mathcal{P}_{\mathcal{R}}$ are from the abundance of lightest supersymmetric particles produced by PBHs radiation and the stable Planck-mass relics after its evaporation.  The light PBHs with mass $M_{\rm pbh}\lesssim 10^{9}~{}\rm g$ are free from BBN (nuclear reactions) constraints due to the shorter lifetime ($\tau\lesssim 0.01~{}s$). Recently, it has been found that the evaporation of light PBHs with masses $10^{8}\lesssim M_{\rm pbh}\lesssim 10^{9}~{}\rm g$ can modify the expansion rate of the Universe and the baryon-to-photon ratio, leading to the changes in the primordial abundance of light nuclei. Therefore, the observed abundance of light nuclei can be used to get the limits on the initial mass fraction of PBHs.
Due to the Hawking radiation, PBHs with mass $M_{\rm PBH}\lesssim 10^{15}\rm g$ have evaporated completely by now. Recently, the ‘memory burden’ effect has been proposed, which is not considered in Hawking radiation and becomes crucial when a BH evaporate half its mass. The mass loss rate of a memory-burdened BH is slowed down, resulting in the existence of light PBHs by now. Therefore, light PBHs with mass $M_{\rm pbh}\lesssim 10^{15}~{}\rm g$ can still radiate high energy particles by now. Moreover, two memory-burdened PBHs could merge to form a new BH with a mass approximately equal to the sum of the two predecessors, and the new formed BH can emit high energy particles due to the Hawking radiation. The observed high energy particles, e.g., neutrinos and photons, can be used to derive the limits on the mass fraction of these light PBHs, which correspond to the small scales of primordial curvature perturbation.
Based on these recent theoretical research process of BH and the limits on the (initial) mass fraction of light PBHs with masses $10^{4}\lesssim M_{\rm PBH}\lesssim 10^{10}~{}\rm g$, we have derived new constraints on the primordial curvature perturbation power spectrum on small scales $1.5\times 10^{18}\lesssim k\lesssim 2.5\times 10^{21}~{}\rm Mpc^{-1}$, which have rarely been studied in previous literature. The strongest limits are from the memory-burdened PBHs. The future detection of high energy particles, e.g., neutrinos, by the telescopes IceCube-Gen2 and GRAND200k can give stronger constraints Chianese et al. (2024); Aartsen et al. (2019, 2021); Álvarez-Muñiz et al. (2020).
IV Acknowledgements
This work is supported by the Shandong Provincial Natural Science Foundation (Grant No.ZR2021MA021).